OpenStudy (anonymous):

a pc. of wire 42 cm. long is bent into the shape of a rectangle whose width is twice its length. find the dimensions of the rectangle??

OpenStudy (diyadiya):

length = l width = 2l 2 ( l+2l ) = 42 2 (3l) = 42 6l =42 l= 42/6 l=7 width =2x7=14

OpenStudy (saifoo.khan):

2(2x + x) = 42

OpenStudy (saifoo.khan):

Like she did.